The area of a rectangular wall of a barn is 54 square feet. its length is 12 feet longer than twice its width. Find the length and width of the wall of the barn.
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = (12 + 2x)
Area = Width x Length
54 = x(12 + 2x)
54 = 12x + 2x^2
Rearrange
2x^2 + 12x - 54 = 0
(2x - 6)(x + 9) = 0
x + 9 = 0
x = -9 (Discounted as -ve)
2x - 6 = 0
2x = 6
x = 3
Width of wall = 3 ft
Length of wall = 18 ft
Hope this helps :-)